# Describe Where You Are: Improving Noise-Robustness for Speech Emotion Recognition with Text Description of the Environment

### v1

@@ 4.1 Data preparation
Our experiment uses the MSP-Podcast corpus, which consists of natural and diverse emotional speech samples from various podcast recordings [6]. The audios do not include background music or overlapped speech, and their predicted SNR is above 20 dB. We consider this corpus a clean emotion speech database for these reasons. This study focuses on predicting the emotional attributes of arousal (calm to active), dominance (weak to strong), and valence (negative to positive). Labels for these attributes were annotated by at least five raters using a seven-point Likert scale. We average the scores provided by raters for each sample to establish its ground truth values. This paper uses version 1.10 of the corpus, which consists of 104,267 annotated utterances. We use the train set to fine-tune the pre-trained speech representation model, using it as the original SER model. We use samples from the development set to select the best model during the fine-tuning process.
@@
We simulate real-world noisy environments by collecting noise sounds from the Freesound repository [39], which contains publicly available ambient noise sounds. We use diverse queries related to each environment to collect noise sounds, including indoor, outdoor, and in-vehicle conditions. We use 20 noisy environments to contaminate the training and development set, consisting of {mall, restaurant, office, airport, station, city, park, street, traffic, home, kitchen, living room, bathroom, bedroom, metro, bus, car, construction site, pedestrian, beach}. For the evaluation, we use six environments, including {plaza, garden, school, tram, sea, boat}. Although these noise sounds are not used during adaptation, they have common characteristics with the noise sounds used during adaptation (e.g. indoor, outdoor, or in-vehicle conditions). We want to evaluate if our proposed method can capture this semantic similarity during the inference. We randomly pick the noise sounds to contaminate the Test1 set of the clean MSP-Podcast corpus. We repeat this process 10 times, creating 10 different sets for three different SNR levels, 5dB, 0dB, and -5dB.
@@ 4.2 Fine-Tuning Transformer-Based Architecture
We implement our proposed approach with two different pre-trained speech representation models: wav2vec2-large-robust [40] and the wavlm-base-plus models [3]. The wav2vec2-large-robust model has shown good performance in the emotional attribute prediction task [32]. The wavlm-base-plus model has shown good performance for emotion recognition in the speech processing universal performance benchmark (SUPERB) [41]. This model is pre-trained with noise, creating representations that are expected to be more robust to noise than other SSL representations. We fine-tune the transformer encoder of the pre-trained speech representation model and the downstream head with the clean version of the MSP-Podcast corpus. For wav2vec2-large-robust, we remove the top 12 transformer layers from the model to preserve the recognition performance with fewer parameters [32]. We import the pre-trained models from the HuggingFace library [42]. We use two fully connected layers for the downstream head, where each layer has 512 nodes, layer normalization, and the rectified linear unit (ReLU) as the activation function. We use dropout in all the hidden layers to increase regularization, with a rate set to $p=0.5$. We use a linear output layer with three nodes to predict emotional attribute scores, where each node predicts the scores for arousal, dominance, and valence. We apply average pooling on top of the last transformer layer’s representation to feed it to the downstream head.
During fine-tuning, we apply Z-normalization to the raw waveform by using the mean and standard deviation estimated over the training set and min-max normalization to the emotional labels, mapping them to the range of 0 to 1. We use 32 utterances per mini-batch and update the model for ten epochs. We use the Adam optimizer [43] with a learning rate warmup scheduling, which shows good performance when fine-tuning a pre-trained transformer architecture [44]. For the first 1,000 mini-batches, we linearly increase the learning rate from $1e^{-8}$ to $1e^{-5}$. After the 1,000 mini-batches, we fix the learning rate to $1e^{-5}$. All of our experiments are conducted on a single NVIDIA GeForce RTX 3090.
@@ 4.4 Baselines
Retrain the original model with noisy speech (RT): This baseline updates the transformer encoder and the downstream head of the Original model with noisy speech. It does not use environmental information during adaptation and inference. As described in Section 4.1, it uses 20 environmental conditions for adaptation. The evaluation uses six other environmental conditions.
Domain adversarial training (DAT): Inspired by Huang et al. ’s work [45], we test a domain adversarial training strategy to adapt an SER model to multiple noisy conditions. Along with the downstream head for the SER task, we attach an environment classifier on top of the average-pooled transformer representations. The environment classifier has the same architecture as the downstream head for the SER task. The environment classifier is trained to minimize the cross-entropy loss between the predicted and the ground-truth noise types. We applied a gradient reversal layer (GRL) between the environment classifier and the transformer encoder to train the transformer encoder to normalize the environment information in the resulting representations. Like the RT baseline, this baseline does not use environmental information during inference.
Table 1: Average CCC of the ten experiments for the proposed text-guided environment-aware methods and the baselines. We report the performance with models implemented using either the wav2vec2-large-robust or wavlm-base-plus feature vectors. We denote with , , and when a model shows significantly better performance than the Original, RT, and DAT models, respectively. We highlight in bold the best performance per condition.
@@
wav2vec2-large-robust
5dB Original 0.59 0.51 0.40
RT 0.60 0.52 0.45
DAT 0.61 0.5 0.45
TG-EAT-CL 0.62 0.52 0.46
TG-EAT-LLM 0.63 0.53 0.47
0dB Original 0.53 0.47 0.33
RT 0.56 0.46 0.4
DAT 0.54 0.44 0.4
TG-EAT-CL 0.52 0.42 0.4
TG-EAT-LLM 0.57 0.48 0.41
-5dB Original 0.27 0.25 0.14
RT 0.24 0.22 0.19
DAT 0.24 0.22 0.16
TG-EAT-CL 0.21 0.2 0.18
TG-EAT-LLM 0.29 0.27 0.21
wavlm-base-plus
5dB Original 0.53 0.46 0.45
RT 0.57 0.48 0.41
DAT 0.59 0.49 0.48
TG-EAT-CL 0.57 0.47 0.47
TG-EAT-LLM 0.58 0.48 0.45
0dB Original 0.40 0.32 0.35
RT 0.53 0.43 0.34
DAT 0.53 0.45 0.42
TG-EAT-CL 0.52 0.43 0.42
TG-EAT-LLM 0.55 0.45 0.40
-5dB Original 0.11 0.07 0.12
RT 0.19 0.12 0.13
DAT 0.22 0.17 0.21
TG-EAT-CL 0.18 0.12 0.19
TG-EAT-LLM 0.29 0.21 0.23
@@ 5.1 Emotion recognition performance
We report the SER performance of our text-guided environment-aware training with our baselines. As described in Section 4.1, we use ten different evaluation sets for three SNR levels. We report the average CCC of ten experiments for each SNR level. We conduct a one-tailed Welch’s t-test between the baselines and our proposed models to assess if the training strategy shows significantly better SER performance in noisy conditions. We assert significance at $p$-value $<$ 0.05.
@@
Table 1 illustrates the SER performance of each model in noisy testing environments. When comparing our baselines (RT, DAT) with the original model, they do not consistently yield performance improvement for all the attributes. RT does not improve the performance neither for arousal and dominance with the wav2vec2-large-robust feature vector, nor for valence with the wavlm-base-plus feature vector. Although the DAT shows significant performance improvement with wavlm-base-plus feature vector, it fails to improve arousal and dominance prediction performance with wav2vec2-large-robust feature vector. Since these baselines do not use environmental information, we can observe the importance of using environmental information when adapting the SER model to multiple noisy environments.
@@
Compared with the baselines, our proposed TG-EAT-LLM performs best when using the wav2vec2-large-robust feature vector. In the 5dB condition, TG-EAT-LLM improves the original model’s performance by 6.7 % (arousal), 3.9% (dominance), and 17.5% (valence). It does not yield the best performance with the wavlm-base-plus feature vector in the 5dB conditions. However, as the SNR level decreases, TG-EAT-LLM shows higher performance than the baselines. In -5dB condition, TG-EAT-LLM shows performance gains of 31.8% (arousal), 23.5% (dominance), and 9.5% (valence) compared to the best baseline, DAT. In spite of having a mismatch in SNR and environment conditions, TG-EAT-LLM shows robust results under all the conditions. These results indicate that guiding the SER model with LLM-based representation can improve the noise-robustness for the SER task. It shows good generalization to unknown environments. Although the DAT approach is effective when using the wavlm-base-plus model for noise conditions above 0dB SNR, using LLM-based representation is more helpful when dealing with low SNR conditions.
@@
When we compare the TG-EAT-CL and TG-EAT-LLM models, we conclude that the CL-based representation does not show a performance improvement over the original SER model, especially with the wav2vec2-large-robust feature vector. We can clearly see that the TG-EAT-CL model does not improve the performance for arousal and dominance in the 0dB and -5dB conditions. This result indicates that pre-training the text encoder to have enriched semantic information is more helpful for the noise-robust SER model than pre-training the text encoder with a audio-text pair.
@@
Figure 2: Embedding differences in the first and the last transformer encoder layers using clean and noisy speech in the -5dB condition. We use the wavlm-base-plus feature vector in this analysis. (a) it extracts and compares clean and noisy representations from each of the models. (b) it extracts and compares the clean representation from the Original model and the noisy representation from each of the models.
@@ 5.2 Embedding analysis
Section 5.1 demonstrated that the TG-EAT-LLM approach shows better performance than the environment-agnostic baselines and the TG-EAT-CL approach. Our initial assumption is that the proposed TG-EAT-LLM can learn appropriate denoising functions for the transformer encoder. To verify this assumption, we analyze the difference between the clean and noisy representations (Fig. 2(a)). We use the wavlm-base-plus feature vector and the noisy speech from the -5dB condition for this analysis. The first analysis compares the clean and noisy representation extracted from each model. We want to assess with this analysis if the model has robustness between clean and noisy speech. The second analysis compares the clean representation from the Original framework and the noisy representation from each of the models (Fig. 2(b)). In this analysis, we want to assess if the model can keep the knowledge of the original SER model. We extract the representations from the first and the last transformer encoder layers and then calculate the mean square difference between clean and noisy representations for each layer.
@@
Compared with the TG-EAT-LLM method, TG-EAT-CL shows a higher representation difference in the first layer. When comparing the clean and noisy representations from the same model, TG-EAT-LLM shows 7.7% less representation difference than the TG-EAT-CL method in the first transformer layer. However, TG-EAT-CL shows less representation difference than the TG-EAT-LLM in the last layer. Even though the downstream head uses the representation from the last transformer layer, TG-EAT-CL shows worse performance than the TG-EAT-LLM approach. LLM-based representation can better denoise the acoustic representation than the CL-based representation. In addition, we speculate that the embedding difference in the lower transformer layer might be the crucial factor for increasing the noise-robustness of the SER system.
(a) TG-EAT-CL
(b) TG-EAT-LLM
@@
We also investigate if the proposed text-based environment embedding clusters similar environments together, which it the key premise of the proposed approach to deal with unseen environments. First, we extract the 26 different text embeddings using the different templates to describe each environmental condition. We project these embeddings into the 2D space to visualize the embedding space using the uniform manifold approximation and projection (UMAP) method [46]. Figure 3 illustrates the text embedding space of TG-EAT-CL and TG-EAT-LLM. The figure shows that both frameworks cluster environmental conditions that are semantically similar together. For example, we observe the embeddings for "boat" and “sea,” together. We also observe the ones for "subway" and "station" clustered together. Both encoders cluster the house environments (house, home, kitchen) and the vehicle environments (bus, taxi, car), which indicates that the text encoder can cluster acoustically similar environments. This analysis implies that our proposed frameworks can deal with unseen environments by clustering acoustically and semantically similar environments.
@@
Table 3: Average CCC of the ten experiments for the unseen environment. We compare the proposed method with the baselines by using the wavlm-base-plus model.
@@
AST 0.60 0.49 0.48
TG-EAT-LLM 0.58 0.48 0.45
0dB Original 0.40 0.32 0.35
RT 0.53 0.43 0.34
GloVe 0.53 0.43 0.37
AST 0.56 0.46 0.43
TG-EAT-LLM 0.55 0.45 0.40
-5dB Original 0.11 0.07 0.12
RT 0.19 0.12 0.13
GloVe 0.25 0.17 0.20
AST 0.25 0.18 0.21
TG-EAT-LLM 0.29 0.21 0.23

### v2

@@ 4.1 Data preparation
Our experiment uses the MSP-Podcast corpus [Busso_2025], which consists of natural and diverse emotional speech samples from various podcast recordings [Lotfian_2019_3]. The audios do not include background music or overlapped speech, and their predicted SNR is above 20 dB. We consider this corpus a clean emotion speech database for these reasons. This study focuses on predicting the emotional attributes of arousal (calm to active), dominance (weak to strong), and valence (negative to positive). Labels for these attributes were annotated by at least five raters using a seven-point Likert scale. We average the scores provided by raters for each sample to establish its ground truth values. This paper uses version 1.10 of the corpus, which consists of 104,267 annotated utterances. We use the train set to fine-tune the pre-trained speech representation model, using it as the original SER model. We use samples from the development set to select the best model during the fine-tuning process.
Table 1: Keywords that are used for contaminating training, development, and testing sets. Freesound illustrates the keywords that are used for crawling the ambient recordings from the Freesound repository. DEMAND illustrates the keywords paired with the recorded sounds in the DEMAND corpus.
Data Split Corpus Keywords
Training, Development Freesound mall, restaurant, office, airport, station, city, park, street, traffic, home, kitchen, living room, bathroom, bedroom, metro, bus, car, construction site, pedestrian, beach
Testing Freesound plaza, garden, school, tram, sea, boat, amusement park, aquarium, arcade, art gallery, backyard, balcony, bank, bar, barn, beach, bridge, cafe, campground, canyon, carnival, cave, cemetery, church, circus, classroom, creek, crowd, dessert, dock, elevator, exhibition hall, factory, fairground, farmyard, festival, field, forest, fountain, gallery, gas station, grocery store, gym, harbor, highway, hospital, hotel, ice rink, industrial site, jungle, lake, laundromat, library, lobby, machine shop, market, meadow, mountain, museum, nightclub, parade, parking lot, patio, pet store, playground, pub, river, rooftop, shopping center, stadium, subway, swimming pool, theater, valley, waiting room, warehouse, waterfall, wetland, workshop, yard
DEMAND washroom, kitchen, living room, sports field, river, park, office, hallway, meeting, subway station, cafeteria, restaurant, traffic intersection, town square, cafe terrace, subway, bus, car
@@
We simulate real-world noisy environments by collecting noise sounds from the Freesound repository [Fonseca_2017], which contains publicly available ambient noise sounds. We use diverse queries related to each environment to collect noise sounds, including indoor, outdoor, and in-vehicle conditions. Additionally, we included the DEMAND dataset for additional testing conditions. DEMAND contains 15 different recording conditions that simulate indoor, outdoor, and in-vehicle environments. We directly use the metadata of each recording sample to define the keyword for the testing conditions. Table 1 illustrates the keywords that are used to contaminate train, development, and test sets. We use 20 noisy environments for the train and development sets and 89 environments to contaminate the test set. Although these noise sounds are not used during adaptation, they have common characteristics with the noise sounds used during adaptation (e.g., indoor, outdoor, or in-vehicle conditions). We want to evaluate whether our proposed method can capture this semantic similarity during inference. We randomly pick the noise sounds to contaminate the Test1 set of the clean MSP-Podcast corpus. We repeat this process 10 times, creating 10 different sets for three different SNR levels, 5dB, 0dB, and -5dB. We also create a random set, where the SNR levels are randomly selected from 5dB to -5dB. This set simulates the testing condition in a real-world application where the SNR level varies.
@@ 4.2 Fine-Tuning Transformer-Based Architecture
We implement our proposed approach with two different pre-trained speech representation models: wav2vec2-large-robust [Hsu_2021_2] and the wavlm-base-plus models [Chen_2022]. The wav2vec2-large-robust model has shown good performance in the emotional attribute prediction task [Wagner_2023]. The wavlm-base-plus model has shown good performance for emotion recognition in the speech processing universal performance benchmark (SUPERB) [Yang_2021_3]. This model is pre-trained with noise, creating representations that are expected to be more robust to noise than other SSL representations. We fine-tune the transformer encoder of the pre-trained speech representation model and the downstream head with the clean version of the MSP-Podcast corpus. For wav2vec2-large-robust, we remove the top 12 transformer layers from the model to preserve the recognition performance with fewer parameters [Wagner_2023]. We import the pre-trained models from the HuggingFace library [Wolf_2019]. We use two fully connected layers for the downstream head, where each layer has 512 nodes, layer normalization, and the rectified linear unit (ReLU) as the activation function. We use dropout in all the hidden layers to increase regularization, with a rate set to $p=0.5$. We use a linear output layer with three nodes to predict emotional attribute scores, where each node predicts the scores for arousal, dominance, and valence. We apply average pooling on top of the last transformer layer’s representation to feed it to the downstream head.
During fine-tuning, we apply Z-normalization to the raw waveform by using the mean and standard deviation estimated over the training set and min-max normalization to the emotional labels, mapping them to the range of 0 to 1. We use the same mean and standard deviation estimated over the training set to normalize the test set’s raw waveform. We use 32 utterances per mini-batch and update the model for ten epochs. We use the Adam optimizer [Kingma_2014_2] with a learning rate warmup scheduling, which shows good performance when fine-tuning a pre-trained transformer architecture [Popel_2018]. For the first 1,000 mini-batches, we linearly increase the learning rate from $1e^{-8}$ to $1e^{-5}$. After the 1,000 mini-batches, we fix the learning rate to $1e^{-5}$. All of our experiments are conducted on a single NVIDIA GeForce RTX 3090.
@@ 4.4 Baselines
Retrain the original model with noisy speech (RT): This baseline updates the transformer encoder and the downstream head of the Original model with noisy speech. It does not use environmental information during adaptation and inference. As described in Section 4.1, it uses 20 environmental conditions for adaptation. The evaluation uses 89 other environmental conditions.
Domain adversarial training (DAT): Inspired by Huang et al. [Huang_2022], we test a domain adversarial training strategy to adapt an SER model to multiple noisy conditions. Along with the downstream head for the SER task, we attach an environment classifier on top of the average-pooled transformer representations. The environment classifier has the same architecture as the downstream head for the SER task. The environment classifier is trained to minimize the cross-entropy loss between the predicted and the ground-truth noise types. We applied a gradient reversal layer (GRL) between the environment classifier and the transformer encoder to train the transformer encoder to normalize the environment information in the resulting representations. Like the RT baseline, this baseline does not use environmental information during inference.
Enhance the noisy speech (SE): This baseline denoises the input noisy speech before feeding it into the original SER model. We use the frequency recurrent convolutional recurrent network (FRCRN) framework [Zhao_2022] to enhance the input speech. The FRCRN model is trained with the 4th DNS challenge dataset, achieving one of the top performances in this challenge [Dubey_2022].
Table 2: Average CCC for models using wav2vec2-large-robust feature vectors. We denote with , , , and when a model shows significantly better performance than the Original, RT, DAT, and SE models, respectively. We also mark and when a baseline significantly perform better than the TG-EAT-CL and TG-EAT-LLM, respectively. We highlight in bold the best performance per condition.
@@
Clean
Original ($\ast$}) 0.63 0.53 0.41
RT ($\dagger$}}) 0.63 0.53 0.46
DAT ($\star$}) 0.63 0.51 0.45
SE ($\ddagger$}}) 0.53 0.48 0.37
TG-EAT-CL ($\diamondsuit$}) 0.63 0.53 0.45
TG-EAT-LLM ($\clubsuit$}) 0.63 0.53 0.46
5dB
Original ($\ast$}) 0.60 0.51 0.40
RT ($\dagger$}}) 0.63 0.52 0.44
DAT ($\star$}) 0.62 0.50 0.44
SE ($\ddagger$}}) 0.50 0.44 0.35
TG-EAT-CL ($\diamondsuit$}) 0.62 0.51 0.45
TG-EAT-LLM ($\clubsuit$}) 0.62 0.52 0.46
0dB
Original ($\ast$}) 0.54 0.46 0.31
RT ($\dagger$}}) 0.55 0.46 0.38
DAT ($\star$}) 0.54 0.44 0.39
SE ($\ddagger$}}) 0.47 0.41 0.35
TG-EAT-CL ($\diamondsuit$}) 0.52 0.42 0.38
TG-EAT-LLM ($\clubsuit$}) 0.56 0.47 0.39
-5dB
Original ($\ast$}) 0.26 0.24 0.11
RT ($\dagger$}}) 0.22 0.21 0.15
DAT ($\star$}) 0.24 0.22 0.13
SE ($\ddagger$}}) 0.23 0.19 0.19
TG-EAT-CL ($\diamondsuit$}) 0.21 0.20 0.15
TG-EAT-LLM ($\clubsuit$}) 0.28 0.26 0.16
Random
Original ($\ast$}) 0.47 0.41 0.29
RT ($\dagger$}}) 0.47 0.39 0.35
DAT ($\star$}) 0.47 0.39 0.34
SE ($\ddagger$}}) 0.37 0.32 0.30
TG-EAT-CL ($\diamondsuit$}) 0.44 0.37 0.34
TG-EAT-LLM ($\clubsuit$}) 0.50 0.42 0.36
Table 3: Average CCC for models using wavlm-base-plus feature vectors. We use the same notations as in Table 2. We highlight in bold the best performance per condition.
SNR Model Arousal Dominance Valence
Clean
Original ($\ast$}) 0.60 0.49 0.46
RT ($\dagger$}}) 0.59 0.49 0.43
DAT ($\star$}) 0.58 0.48 0.48
SE ($\ddagger$}}) 0.58 0.46 0.43
TG-EAT-CL ($\diamondsuit$}) 0.57 0.47 0.47
TG-EAT-LLM ($\clubsuit$}) 0.59 0.48 0.46
5dB
Original ($\ast$}) 0.54 0.45 0.44
RT ($\dagger$}}) 0.58 0.48 0.41
DAT ($\star$}) 0.58 0.48 0.47
SE ($\ddagger$}}) 0.55 0.45 0.40
TG-EAT-CL ($\diamondsuit$}) 0.57 0.47 0.46
TG-EAT-LLM ($\clubsuit$}) 0.58 0.47 0.44
0dB
Original ($\ast$}) 0.40 0.31 0.33
RT ($\dagger$}}) 0.53 0.43 0.33
DAT ($\star$}) 0.53 0.45 0.41
SE ($\ddagger$}}) 0.53 0.44 0.41
TG-EAT-CL ($\diamondsuit$}) 0.51 0.42 0.40
TG-EAT-LLM ($\clubsuit$}) 0.55 0.45 0.38
-5dB
Original ($\ast$}) 0.11 0.07 0.10
RT ($\dagger$}}) 0.18 0.11 0.12
DAT ($\star$}) 0.22 0.16 0.17
SE ($\ddagger$}}) 0.28 0.22 0.23
TG-EAT-CL ($\diamondsuit$}) 0.17 0.11 0.18
TG-EAT-LLM ($\clubsuit$}) 0.29 0.20 0.20
Random
Original ($\ast$}) 0.34 0.25 0.31
RT ($\dagger$}}) 0.45 0.33 0.30
DAT ($\star$}) 0.46 0.37 0.38
SE ($\ddagger$}}) 0.44 0.35 0.35
TG-EAT-CL ($\diamondsuit$}) 0.43 0.33 0.37
TG-EAT-LLM ($\clubsuit$}) 0.50 0.39 0.36
@@ 5.1 Emotion recognition performance
We report the SER performance of our text-guided environment-aware training with our baselines. As described in Section 4.1, we use ten different evaluation sets for three SNR levels. We report the average CCC of ten experiments for each SNR level. We conduct a one-tailed Welch’s t-test between the baselines and our proposed models to assess if the training strategy shows significantly better SER performance in noisy conditions. We assert significance at $p$-value << 0.05.
@@
Tables 2 and 3 illustrate the SER performance of each model in noisy testing environments. When comparing our baselines (RT, DAT, SE) with the original model, they do not consistently yield performance improvement for all the attributes. RT does not improve performance for either arousal or dominance with the wav2vec2-large-robust feature vector, or for valence with the wavlm-base-plus feature vector. Although the DAT and SE show significant performance improvements with the wavlm-base-plus feature vector, both baselines fail to improve arousal and dominance prediction performance with the wav2vec2-large-robust feature vector. Since these baselines do not use environmental information, we can observe the importance of incorporating it when adapting the SER model to multiple noisy environments.
@@
Compared with the baselines, our proposed TG-EAT-LLM performs the best when using the wav2vec2-large-robust feature vector. In the random condition, TG-EAT-LLM improves the original model’s performance by 6.3% (arousal), 2.4% (dominance), and 24.1% (valence). It yields the best performance with the wavlm-base-plus feature vector for arousal and dominance prediction tasks. In the random condition, TG-EAT-LLM shows performance gains of 8.6% (arousal) and 5.4% (dominance) compared to the best baseline, DAT. Unlike with the wav2vec2-large-robust representation, DAT significantly improves the original model’s performance for all the attributes with the wavlm-base-plus representation. The wavlm-base-plus is pre-trained with noisy data, while the wav2vec2-large-robust is trained with a diverse speech corpus under clean conditions. This difference makes the wavlm-base-plus inherently more robust to noise, which leads to the successful improvement with the baselines that do not use environmental information. We note that TG-EAT-LLM consistently outperforms the original model across all SSL representations. These results indicate that guiding the SER model with LLM-based representation can improve the noise-robustness for the SER task. It shows good generalization to unknown environments.
For the valence prediction, the SE baseline shows the best performance under the -5dB condition. Previous studies have shown that valence performance correlates with the linguistic information [Wagner_2023]. This phenomenon could explain how the SE baseline can improve valence performance by explicitly enhancing speech intelligibility. However, it does not always yield the best performance for arousal and dominance. Both arousal and dominance are related to acoustic characteristics rather than linguistic information. Thus, this observation implies that the enhancement module can manipulate the acoustic characteristics of the original speech. We can also see that the SE baseline does not yield the best performance for valence under 5dB conditions, where the impact of acoustic distortion could be higher than the impact of intelligibility improvement.
@@
When we compare the TG-EAT-CL and TG-EAT-LLM models, we conclude that the CL-based representation does not show a performance improvement over the original SER model, especially with the wav2vec2-large-robust feature vector. We can clearly see that the TG-EAT-CL model does not improve the performance for arousal and dominance in the 0dB and -5dB conditions. This result indicates that pre-training the text encoder to have enriched semantic information is more helpful for the noise-robust SER model than pre-training the text encoder with an audio-text pair.
An interesting and counter-intuitive finding here is that the models trained with noisy speech (e.g., RT, DAT, TG-EAT-CL, TG-EAT-LLM) outperform the Original model under clean conditions when experimenting with the wav2vec2-large-robust architecture. We assume this improvement is caused by exposing the noisy speech to the wav2vec2-large-robust representation, which is not trained with noisy speech in its pre-training stage. Previous studies have shown that augmenting the training set with multiple conditions not only improves automatic speech recognition (ASR) performance under noisy conditions but also improves under a clean condition [Ko_2017, Pesoparada_2022]. As discussed in the previous section, we hypothesize that improvements in speech intelligibility leads to improvement of valence prediction. Based on these observations, we conclude that this phenomenon, while unintuitive, demonstrates the benefits of data augmentation under clean conditions.
@@
Figure 2: Embedding differences in the first and the last transformer encoder layers using clean and noisy speech in the -5dB condition. We use the wavlm-base-plus feature vector in this analysis. (a) illustrates the mean square error (MSE) between the clean and noisy representations, where both representations are extracted from each of the final models. (b) illustrates the MSE between the clean representation extracted from the Original model and the noisy representation extracted from each final model.
@@ 5.2 Embedding analysis
Section 5.1 demonstrated that the TG-EAT-LLM approach shows better performance than the environment-agnostic baselines and the TG-EAT-CL approach. Our initial assumption is that the proposed TG-EAT-LLM can learn appropriate denoising functions for the transformer encoder. To verify this assumption, we analyze the difference between the clean and noisy representations (Fig. 2(a)). We use the wavlm-base-plus feature vector and the noisy speech from the -5dB condition for this analysis. The first analysis compares the clean and noisy representation extracted from each model. We want to assess with this analysis if the model is robust by comparing the representation obtained with clean and noisy speech. The second analysis compares the clean representation from the Original framework and the noisy representation from each of the models (Fig. 2(b)). In this analysis, we want to assess if the model can keep the knowledge of the original SER model. We extract the representations from the first and the last transformer encoder layers and then calculate the mean square difference between clean and noisy representations for each layer.
@@
Compared with the TG-EAT-LLM method, TG-EAT-CL shows a higher representation difference in the first layer. When comparing the clean and noisy representations from the same model, TG-EAT-LLM shows 7.7% less representation difference than the TG-EAT-CL method in the first transformer layer. However, TG-EAT-CL shows less representation difference than the TG-EAT-LLM in the last layer. Even though the downstream head uses the representation from the last transformer layer, TG-EAT-CL shows worse performance than the TG-EAT-LLM approach. LLM-based representation can better denoise the acoustic representation than the CL-based representation. In addition, we speculate that the embedding difference in the lower transformer layer might be the crucial factor for increasing the robustness to noise of the SER system.
((a))
((b))
@@
We also investigate if the proposed text-based environment embedding clusters similar environments together, which is the key premise of the proposed approach to deal with unseen environments. First, we randomly select 21 different keywords, each representing an indoor, outdoor, and in-vehicle environment. Each environment includes seven keywords extracted from the train and test sets, aiming to illustrate the model’s capability to cluster similar environments in both seen and unseen environments. We extract the text embedding from these 21 keywords by using the same template that we used for our TG-EAT frameworks (i.e., “This speech is recorded in {environment}.”) We project these embeddings into the 2D space to visualize the embedding space using the uniform manifold approximation and projection (UMAP) method [Mcinnes_2018]. Figure 3 illustrates the text embedding space of TG-EAT-CL and TG-EAT-LLM. The figure shows that both frameworks cluster semantically similar environmental conditions together. For example, we observe the embeddings for “boat” and “sea,” together. We also observe the ones for “subway” and “station” clustered together. Both encoders cluster the house environments (“house”, “home”, “kitchen”) and the vehicle environments (“bus”, “taxi”, “car”), which indicates that the text encoder can cluster acoustically similar environments. This analysis implies that our proposed frameworks can handle unseen environments by clustering acoustically and semantically similar environments.
Table 4: Silhouette score of text embedding space of TG-EAT-CL and TG-EAT-LLM. We apply K-means clustering on the projected environmental embedding with $K$ clusters (K=3,5,7K={3,5,7}).
K = 3 K = 5 K = 7
TG-EAT-CL 0.11 0.10 0.10
TG-EAT-LLM 0.57 0.43 0.40
To provide a quantitative analysis of the environmental embeddings and their impact on the model’s representations, we evaluated the clustering quality of the environmental text embeddings. We extracted embeddings for all environments listed in Table 1 from both the TG-EAT-CL and TG-EAT-LLM text encoders. We calculated the silhouette score for each set of embeddings using the K-means clustering [Shahapure_2020]. Table 4 illustrates the silhouette score of each embedding projection with a different number of clusters. With three clusters, the TG-EAT-LLM embeddings achieved a score of 0.57, substantially higher than the 0.10 score from the TG-EAT-CL embeddings. This result indicates that the LLM-based encoder generates more separable and well-defined clusters for different environments. This higher-quality embedding structure correlates with the superior performance of the TG-EAT-LLM model in noisy conditions, suggesting that more discriminable environmental representations are key to achieving robust performance.
@@
Table 6: Average CCC of the ten experiments for the unseen environment. We compare the proposed method with the baselines by using the wavlm-base-plus model. We denote with when a model shows significantly better performance than the Original model.
@@
AST 0.59 0.49 0.41
TG-EAT-LLM 0.58 0.48 0.44
0dB
Original 0.40 0.31 0.33
RT 0.53 0.43 0.33
GloVe 0.53 0.42 0.37
AST 0.55 0.44 0.34
TG-EAT-LLM 0.55 0.45 0.38
-5dB Original 0.11 0.07 0.10
RT 0.18 0.11 0.12
GloVe 0.24 0.16 0.18
AST 0.28 0.20 0.14
TG-EAT-LLM 0.29 0.20 0.20